Fermi National Laboratory


Accelerator Update

Monday March 8
The day shift began with Operations monitoring store 3277, with Antiproton Source (Pbar) stacking, and with Main Injector (MI), Recycler, and MiniBooNE experts conducting studies.

At 11:13 AM, the vacuum technicians accessing the Switchyard (SY) F1 Manhole inadvertently dropped the interlocks. Operators re-interlocked the area and the techs finished their work on the SEPTA (FSEP).

A MI kicker power supply (I:KPS4A) tripped off at 2:37 PM. Experts found a tripped breaker. They had the supply operating normally by 3:05 PM. Pbar resumed stacking a minute later.

Operations reported at 4:30 PM that the MI/SY slow acceleration studies were occasionally causing all the MI RF stations to trip off. This trip would halt stacking until the stations got turned back on.

All MI studies had finished by 5:55 PM, but then started up again 8:30 PM and ran for another hour.

Tevatron (TeV) experts performed end-of-store studies at 10 PM and then Operations terminated store 3277 at 11:38 PM.

Operations put the TeV through a dry squeeze at 11:53 PM.

Tuesday March 9
The midnight shift began with the TeV, MI, and Pbar preparing for shot setup.

Operations reported at 12:15 AM that they had been investigating a Linac RF station (LRF2) trip that occurred around midnight. Operators called a Linac tech and determined that the station was "hosed," a technical term that will get explained later. The tech came in to work on the problem. With this station off, it put shot setup on hold. At 4:10 AM, the tech found what he thought was a bad relay, but was having trouble finding a replacement.

At 4:52 AM, Fermilab suffered a power glitch. Most devices tripped off, but amazingly the TeV didn't quench and we still maintained the stack and the stash. The Duty Electrician stated that feeder 35 faulted.

CHL reported at 5 AM that they lost a helium compressor, but their cryo systems were relatively stable. Some of the wet engines also tripped off.

At 8:55 AM the Tev cryo recovery was complete, however some networking problems still existed.

At 11:30 AM Linac decided to switch to the I- source because of problems with the H- source. Change out was complete by noon.

Stacking and beam to MiniBooNE resumes at 12:08 PM.

The Tev was ramping by 1:30 PM.

Beam was injected into the Tev at 3:15 PM.

Shot setup for the Tev began at 7:00 PM.

Problems with Linac delayed shot setup until 9:00 PM.

Store #3283 was colliding by 10:41 PM with an initial luminosity of 60.95 E30. Pbar then returned to stacking.

Wednesday March 10 Transfers of Pbars to the Recycler started at 2:12 AM and were completed by 2:25 AM.

Today's Plans
Keep the Tev store in until tomorrow day shift.

Accelerator Update Archive

More Information
For Tevatron luminosity charts and the current status of Fermilab's accelerators and detectors (live!), please go to Fermilab Now

Comments and Suggestions
What do you think about the Accelerator Updates? Please send comments and suggestions to: accelupdates@fnal.gov.



last modified 3/10/2004   email Fermilab